OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (anonymous):

put the equation they give you into slope intercept. y= 1/4x+6/4 thats just a line going diagonal, so you make another equation with the same slope which is 1/4 remember y=mx+b and you find the point (-8,5) on a graph and you draw a line from there with a slope of 1/4. and there you go. b= the point of y when x=0 m= slope SOURCE: http://answers.yahoo.com/question/index?qid=20111024211331AAfOVzL

OpenStudy (anonymous):

that confused me .

OpenStudy (anonymous):

for your line, x + 4y = 6 what is the slope?

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

ok, we need to solve for y to get your equation in slope intercept form:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):